# SINDI: an Efficient Index for Approximate Maximum Inner Product Search on Sparse Vectors

### v1

@@ Abstract.
Sparse vector Maximum Inner Product Search (MIPS) is crucial in multi-path retrieval for Retrieval-Augmented Generation (RAG). Recent inverted index-based and graph-based algorithms have achieved high search accuracy with practical efficiency. However, their performance in production environments is often limited by redundant distance computations and frequent random memory accesses. Furthermore, the compressed storage format of sparse vectors hinders the use of SIMD acceleration. In this paper, we propose the sparse inverted non-redundant distance index (Sindi), which incorporates three key optimizations: (i) Efficient Inner Product Computation: Sindi leverages SIMD acceleration and eliminates redundant identifier lookups, enabling batched inner product computation; (ii) Memory-Friendly Design: Sindi replaces random memory accesses to original vectors with sequential accesses to inverted lists, substantially reducing memory-bound latency. (iii) Vector Pruning: Sindi retains only the high-magnitude non-zero entries of vectors, improving query throughput while maintaining accuracy. We evaluate Sindi on multiple real-world datasets. Experimental results show that Sindi achieves state-of-the-art performance across datasets of varying scales, languages, and models. On the MsMarco dataset, when Recall@50 exceeds 99%, Sindi delivers single-thread query-per-second (QPS) improvements ranging from $4.2\times$ to $26.4\times$ compared with Seismic and PyANNs. Notably, Sindi has been integrated into Ant Group’s open-source vector search library, VSAG.
@@ 1. Introduction
Recently, retrieval-augmented generation (RAG) (gao2023retrieval, ; lewis2020retrieval, ; ma2025lightretriever, ) become one of most successful information retrieval framework attracting attention from research communities and industry. Usually, texts are embedded into dense vectors (i.e., no dimension of the vector is zero entry) in RAG, then retrieved through approximate nearest neighbor search (ANNS) on their corresponding dense vectors.
To enhance the RAG framework, researchers find that using sparse vectors retrieval to complement dense vector based RAG can yield better overall accuracy and recall performance. Different from dense vectors, sparse vectors (i.e., only a very small portion of dimensions of spare vectors are non-zero entries) are generated by specific models (e.g., Splade (distill-splade, ; formal2021splade, ; splade, )) to preserve semantic information while enabling precise lexical matching (ma2025lightretriever, ). In the enhanced RAG framework, dense vectors capture the holistic semantic similarity between texts and sparse vectors ensure exact term recall, therefore resulting in better overall performance. We show the process of the enhanced RAG in the following example:
@@
The similarity retrieval problem for sparse vectors is formally known as the Maximum Inner Product Search (MIPS) (ip-1, ; ip-2, ; ip-3, ; ip-4, ; ip-5, ), which aims to identify the top-$k$ vectors in a dataset that have the largest inner product value with a given query vector. However, due to the curse of dimensionality (indyk1998approximate, ), performing exact MIPS in high-dimensional spaces is computationally prohibitive. To mitigate this issue, we focus on Approximate Maximum Inner Product Search (AMIPS), which trades a small amount of recall for significantly improved search efficiency.
Many algorithms (seismic, ; sosia, ; bmp, ) have been proposed for AMIPS, employing techniques such as inverted index, proximity graphs, and hash-based partitioning. They improve efficiency by grouping similar vectors into the same partition, thereby reducing the number of candidates examined during query processing.
@@ 2.1. Problem Definition
For small datasets, exact Sparse-MIPS can be obtained by scanning all vectors. For large-scale high-dimensional collections, this is prohibitively expensive, and Approximate MIPS mitigates the cost by trading a small loss in accuracy for much higher efficiency.
@@ 2.2. Existing Solutions
Representative algorithms for the AMIPS problem on sparse vectors include the inverted-index based Seismic (seismic, ), the graph based PyAnns. Seismic constructs an inverted list based on vector dimensions. PyAnns creates a proximity graph where similar vectors are connected as neighbors.
@@
Redundant Distance Computations. Sparse vectors incur high distance computation cost due to (i) the explicit lookup needed to identify the common dimensions $\Omega(\vec{x},\vec{q})$ between a document $\vec{x}$ and a query $\vec{q}$, resulting in complexity $\mathcal{O}(\lVert q\rVert+\lVert x\rVert)$, and (ii) the inability of existing algorithms to exploit SIMD acceleration for inner product computation. Profiling 6980 queries on the MsMarco dataset (1M vectors) using perf and vtune shows that PyANNs spent 83.3% of CPU cycles on distance calculation.
Random Memory Accesses. The inefficiency of memory access in existing algorithms can be attributed to two main factors. First, they organize similar data points into the same partition and. To improve accuracy, vectors are replicated across multiple partitions. This replication breaks the alignment between storage layout and query traversal order, preventing cache-friendly sequential access. During retrieval, the index returns candidate vector IDs, which incur random memory accesses to fetch their corresponding data, leading to frequent cache misses. Moreover, $\lVert x\rVert$ varies across sparse vectors, requiring offset table lookups to locate each vector’s data. In our measurements, Seismic averaged 5168 random vector accesses per query (5.1 MB), with an L3 cache miss rate of 67.68%.
@@ 3. Full Precision Inverted Index
  • In § 3.1 Sindi constructs a value-based inverted index by storing both vector identifiers and their corresponding dimension values in posting lists. This design eliminates the redundant dimension-matching overhead present in traditional inverted indexes.
  • In § 3.2, Sindi employs a two-phase search process involving product computation and accumulation. By using SIMD instructions in multiplication, it reduces query complexity from $O(\|q\|+\|x\|)$ to $O\Bigl{(}\frac{\|q\|}{s}\Bigr{)}$. This maximizes CPU utilization and improves query throughput.
@@ 3.1. Value-storing Inverted Index
Figure 3 shows the inverted lists constructed for vectors $x_{1}$ to $x_{5}$, comprising five term lists. When a query $q$ arrives, it sequentially probes the inverted lists for dimensions 1, 3, and 5. The right side illustrates the common non-zero dimensions found when computing the inner product between the query and the documents. For example, in dimension 1, the inverted list retrieves $x_{2}$ and $x_{4}$, and the inner-product computation also multiplies $q$ with $x_{2}^{1}$ and $x_{4}^{1}$. Therefore, the document IDs retrieved from the inverted lists overlap exactly with those used in finding common non-zero dimensions for the inner product. This indicates that we can compute the products of these non-zero entries during document retrieval itself.
Inspired by this observation, we extend the inverted list to store not only the ID $i$ of each vector $\vec{x}_{i}$, but also its value $x^{j}$ for the corresponding dimension $j$. This design eliminates the cost of searching $\Omega(\vec{x}_{i},\vec{q})$ during the inner product computation, as well as the random memory access overhead for retrieving $x_{i}$.
@@ 3.2. Efficient Distance Computation
The entire query process of Sindi can be summarized into two stages: (1) Product Computation. Given an incoming query vector $\vec{q}$, for each non-zero $q^{j}$ we fetch the $j$-th inverted list $I_{j}$, compute the products $q^{j}\times x_{i}^{j}$ for all $x_{i}^{j}\in I_{j}$, and temporarily store the results into the array $T^{j}$. (2) Accumulation. The values in $T^{j}$ are accumulated into the distance array $A$. The length of $A$ is set to $\|\mathcal{D}\|$, so that each entry $A[t]$ corresponds uniquely to a vector $\vec{x}_{t}\in\mathcal{D}$, allowing the accumulation from $T^{j}$ to $A$ to be completed in $\mathcal{O}(1)$ time per element. After completing the accumulation for all $T^{j}$ with $q^{j}\in q$, each $A[t]$ contains $A[t]=\delta(\vec{x}_{t},\vec{q})$. If $A[t]=0$, it means that $\Omega(\vec{x}_{t},\vec{q})=\emptyset$.
@@
Figure 4 illustrates a query example. Since $\|\mathcal{D}\|=9$, the distance array $A$ is initialized with $\text{size}(A)=9$ and all elements set to 0. The query $\vec{q}$ contains three non-zero components $q^{1}$, $q^{5}$, and $q^{8}$, thus only the inverted lists $I_{1}$, $I_{5}$, and $I_{8}$ need to be traversed. Take $\vec{x}_{4}$ as an example: in $I_{1}$, the value is $x_{4}^{1}=6.8$, and the product with $q^{1}$is $17.0$, which is temporarily stored in $T[0]$. Accumulating to $A[4]$ gives $A[4]=0+17.0=17.0$. Similarly, in $I_{5}$ we have $x_{4}^{5}\times q^{5}=14.0$, which is stored in $T[2]$; adding $T[2]$ to $A[4]$ yields $A[4]=31.0$. The computation for $I_{8}$ is analogous. Finally, we obtain $A[4]=36.1$, which equals $\delta(\vec{x}_{4},\vec{q})$. Although $\vec{x}_{4}$ has another non-zero entry $x_{4}^{2}$, it does not contribute to the inner product and thus can be ignored. The same accumulation procedure applies to $\vec{x}_{2},\vec{x}_{3},\vec{x}_{6}$, and $\vec{x}_{7}$. Eventually, $A[4]$ has the highest value, so the nearest neighbor of $\vec{q}$ is $\vec{x}_{4}$.
Figure 4. An example of Sindi index and query process.
Using SIMD instructions, Sindi batch-processes the $j$-th inverted list $I_{j}$, multiplying $q^{j}$ with each $x_{i}^{j}$ it contains and writing the results sequentially into $T^{j}$. This not only utilizes CPU resources efficiently, but also reduces the time complexity of the inner product computation from $\mathcal{O}(\|v\|+\|q\|)$ to $\mathcal{O}\!\left(\frac{\|q\|}{s}\right)$, where $s$ is the number of elements processed per SIMD operation. The derivation of the distance computation complexity for Sindi is as follows:
@@
Let $\mathcal{D}$ be the dataset, and let $I$ denote an inverted index, and $I_{j}$ is the $j$-th inverted list, let the set $\mathcal{I}^{j}=\{x_{i}^{j}\mid x_{i}^{j}\in x_{i}\}$ contains the non-zero entries in $I_{j}$.
Given a query vector $\vec{q}$, let $\mathcal{J}=\{j\mid q^{j}\in q\}$ denote the set of dimensions containing non-zero entries in $q$. Let $\mathcal{X}=\{\vec{x}_{i}\mid x_{i}^{j}\in x_{i},j\in\mathcal{J}\}$ be the set of candidate vectors retrieved by $\vec{q}$.
Let $s$ be the number of dimensions that can be processed simultaneously using SIMD instructions. Then, the per-vector time complexity of computing the inner product between $\vec{q}$ and all $\vec{x}_{i}\in\mathcal{X}$ is
$\varTheta\left(\frac{\|q\|}{s}\right).$
Proof. The total number of non-zero entries accessed across all inverted lists corresponding to $\mathcal{J}$ is:
$\sum_{j\in\mathcal{J}}\|\mathcal{I}^{j}\|.$
Since $s$ dimensions can be processed in parallel using SIMD, the total time complexity for computing inner products between $\vec{q}$ and all $\vec{x}_{i}\in\mathcal{X}$ is:
$T_{\text{total}}=\sum_{j\in\mathcal{J}}\frac{\|\mathcal{I}^{j}\|}{s}.$
The amortized complexity per vector is obtained by dividing $T_{\text{total}}$ by the number of candidates $\|\mathcal{X}\|$:
$T=\frac{T_{\text{total}}}{\|\mathcal{X}\|}=\frac{\sum_{j\in\mathcal{J}}\|\mathcal{I}^{j}\|}{s\cdot\|\mathcal{X}\|}.$
In general, for any $\vec{x}_{i}\in\mathcal{X}$, we have $q\cap x_{i}\subseteq q$, implying that $\Omega(=\vec{x}_{i},=\vec{q})$ is at most $\|q\|$. Therefore:
$\sum_{j\in\mathcal{J}}\|\mathcal{I}^{j}\|\leq\sum_{\vec{x}_{i}\in\mathcal{X}}\|q\|.$
Substituting this relation into the expression for $T$ yields:
@@
Hence, the per-vector time complexity is:
$\varTheta\left(\frac{\|q\|}{s}\right).$
@@
In summary, the Sindi index eliminates the overhead of redundant term searches and enables SIMD to compute inner products in batches, fully leveraging the CPU’s computational power. Moreover, it avoids the memory access costs associated with traversing the original vectors. Instead, the required values can be directly retrieved from the lists, allowing in-place product computation.

### v2

# SINDI: An Efficient Index for Sparse Vector Approximate Maximum Inner Product Search

@@ Abstract.
Sparse vector Maximum Inner Product Search (MIPS) is crucial in multi-path retrieval for Retrieval-Augmented Generation (RAG). Recent inverted index-based and graph-based algorithms have achieved high search accuracy with practical efficiency. However, their performance in production environments is often limited by redundant distance computations and frequent random memory accesses. Furthermore, the compressed storage format of sparse vectors hinders the use of SIMD acceleration. In this paper, we propose the sparse inverted non-redundant distance index (Sindi), which incorporates three key optimizations: (i) Efficient Inner Product Computation: Sindi leverages SIMD acceleration and eliminates redundant identifier lookups, enabling batched inner product computation; (ii) Memory-Friendly Design: Sindi replaces random memory accesses to original vectors with sequential accesses to inverted lists, substantially reducing memory-bound latency. (iii) Vector Pruning: Sindi retains only the high-value non-zero entries of vectors, improving query throughput while maintaining accuracy. We evaluate Sindi on multiple real-world datasets. Experimental results show that Sindi achieves state-of-the-art performance across datasets of varying scales, languages, and models. On the MsMarco dataset, when Recall@50 exceeds 99%, Sindi delivers single-thread query-per-second (QPS) improvements ranging from $4.2\times$ to $26.4\times$ compared with Seismic and PyANNS. Notably, Sindi has been integrated into Ant Group’s open-source vector search library, VSAG.
@@
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/Roxanne0321/vsag/tree/sparse.
@@ 1. Introduction
Recently, retrieval-augmented generation (RAG) (Sawarkar et al., 2024; Şakar and Emekci, 2025; Lewis et al., 2020; Su et al., 2025; Quinn et al., 2025; Pogrebinsky et al., 2025) become one of most successful information retrieval framework attracting attention from research communities and industry. Usually, texts are embedded into dense vectors (i.e., no dimension of the vector is zero entry) in RAG, then retrieved through approximate nearest neighbor search (ANNS) on their corresponding dense vectors.
To enhance the RAG framework, researchers find that using sparse vectors retrieval to complement dense vector based RAG can yield better overall accuracy and recall performance. Different from dense vectors, sparse vectors (i.e., only a very small portion of dimensions of spare vectors are non-zero entries) are generated by specific models (e.g., Splade (Formal et al., 2022, 2024; Formal et al., 2021; Lassance and Clinchant, 2022)) to preserve semantic information while enabling precise lexical matching (Ma et al., 2025). In the enhanced RAG framework, dense vectors capture the holistic semantic similarity between texts and sparse vectors ensure exact term recall, therefore resulting in better overall performance. We show the process of the enhanced RAG in the following example:
@@
The similarity retrieval problem for sparse vectors is formally known as the Maximum Inner Product Search (MIPS) (Keivani et al., 2018; Pham, 2021; Song et al., 2021; Yan et al., 2018; Zhao et al., 2023), which aims to identify the top-$k$ vectors in a dataset that have the largest inner product value with a given query vector. However, due to the curse of dimensionality (Indyk and Motwani, 1998), performing exact MIPS in high-dimensional spaces is computationally prohibitive. To mitigate this issue, we focus on Approximate Maximum Inner Product Search (AMIPS) (Song et al., 2021; Abuzaid et al., 2019; Neyshabur and Srebro, 2015), which trades a small amount of recall for significantly improved search efficiency.
Many algorithms (Bruch et al., 2024; Zhao et al., 2024; Mallia et al., 2024; PyANNS Contributors, 2025) have been proposed for AMIPS, employing techniques such as inverted index (Bruch et al., 2024), proximity graphs (PyANNS Contributors, 2025), and hash-based (Zhao et al., 2024) partitioning. They improve efficiency by grouping similar vectors into the same partition, thereby reducing the number of candidates examined during query processing.
@@ 2.1. Problem Definition
For small datasets, exact Sparse-MIPS can be obtained by scanning all vectors. For large-scale high-dimensional collections, this is prohibitively expensive, and Approximate Sparse-MIPS mitigates the cost by trading a small loss in accuracy for much higher efficiency.
@@ 2.2. Existing Solutions
Representative algorithms for the AMIPS problem on sparse vectors include the inverted-index based Seismic (Bruch et al., 2024), the graph based PyANNS (PyANNS Contributors, 2025). Seismic constructs an inverted list based on vector dimensions. PyANNS creates a proximity graph where similar vectors are connected as neighbors.
@@
Redundant Distance Computations. Sparse vectors incur high distance computation cost due to (i) the explicit lookup needed to identify the common dimensions $\Omega(\vec{x},\vec{q})$ between a document $\vec{x}$ and a query $\vec{q}$, resulting in complexity $\mathcal{O}(\lVert q\rVert+\lVert x\rVert)$, and (ii) the inability of existing algorithms to exploit SIMD acceleration for inner product computation. Profiling 6980 queries on the MsMarco dataset (1M vectors) using perf (Linux Kernel Organization, 2025) and VTune (Intel Corporation, 2025) shows that PyANNS spent 83.3% of CPU cycles on distance calculation.
Random Memory Accesses. The inefficiency of memory access in existing algorithms can be attributed to two main factors. First, Seismic organize similar data points into the same partition. To improve accuracy, vectors are replicated across multiple partitions. This replication breaks the alignment between storage layout and query traversal order, preventing cache-friendly sequential access. During retrieval, the index returns candidate vector IDs, which incur random memory accesses to fetch their corresponding data, leading to frequent cache misses. Moreover, $\lVert x\rVert$ varies across sparse vectors, requiring offset table lookups to locate each vector’s data. In our measurements, Seismic averaged 5168 random vector accesses per query (5.1 MB), with an L3 cache miss rate of 67.68%.
@@ 3. Full Precision Inverted Index
  • In § 3.1, Sindi constructs a value-based inverted index by storing both vector identifiers and their corresponding dimension values in posting lists. This design eliminates the redundant dimension-matching overhead present in traditional inverted indexes.
  • In § 3.2, Sindi employs a two-phase search process involving product computation and accumulation. By using SIMD instructions in multiplication, it reduces query complexity from $O(\|q\|+\|x\|)$ to $O\Bigl{(}\frac{\|q\|}{s}\Bigr{)}$. This maximizes CPU utilization and improves query throughput. w
@@ 3.1. Value-storing Inverted Index
Figure 3 illustrates the inverted lists constructed for vectors $x_{1}$ to $x_{5}$. When a query $q$ arrives, it sequentially probes the inverted lists for dimensions 1, 3, and 5. In the base dataset, only the dimensions belonging to $\Omega(\vec{q},\vec{x})$ need to be traversed during the inner product computation. For example, although $x_{4}$ has a value in dimension 2, this dimension is not in $\Omega(\vec{q},x_{4})$ and thus is never accessed during the computation. We further observe that the document identifiers retrieved from the inverted lists overlap exactly with those used to determine the common non-zero dimensions for the inner product. This implies that the products of these non-zero entries can be computed during the document retrieval process itself, thereby ensuring that only dimensions in $\Omega(\vec{q},\vec{x})$ are involved in the inner product computation.
Inspired by these observations, we extend each inverted list $I_{j}$ to store not only the identifier $i$ of the vector $\vec{x}_{i}$, but also the value $x_{i}^{j}$ in dimension $j$. In our notation, we simply write $x_{i}^{j}$ in $I_{j}$ to denote this stored value, with the subscript $i$ implicitly encoding the associated vector ID. This value-storing design eliminates the cost of explicitly locating $\Omega(\vec{x}_{i},\vec{q})$ during inner product computation, as well as the additional random memory access that would otherwise be required to fetch $x_{i}^{j}$ from the original vector.
@@ 3.2. Efficient Distance Computation
With the value-storing inverted index, the inner product between a query $\vec{q}$ and candidate vector $\vec{x}$ can be computed without explicitly identifying $\Omega(\vec{q},\vec{x})$. During search, Sindi organizes the computation into two stages: product computation and accumulation.
In the first stage, as each relevant posting list $I_{j}$ is traversed, the products $q^{j}\times x_{i}^{j}$ are computed (using SIMD instructions when possible) and stored sequentially in a temporary array $T^{j}$. Here, $T^{j}$ holds the products for $I_{j}$, with each entry $T^{j}[t]$ corresponding to the $t$-th entry in $I_{j}$. Therefore, $T^{j}$ has the same length as $I_{j}$.
In the second stage, the values in $T^{j}$ are aggregated into a preallocated distance array $A$ of length $\lVert\mathcal{D}\rVert$, where each entry $A[i]$ stores the accumulated score for vector $\vec{x}_{i}$. This arrangement enables $\mathcal{O}(1)$ time per accumulation into $A$, and naturally exploits SIMD parallelism.
The following example illustrates the detailed steps of the search procedure introduced above.
@@
Figure 4 illustrates the search procedure. Since $\lVert\mathcal{D}\rVert=9$, the distance array $A$ is initialized with $\text{size}(A)=9$ and all elements set to 0. The query $\vec{q}$ contains three non-zero components, $q^{1}$, $q^{5}$, and $q^{8}$, so only the inverted lists $I_{1}$, $I_{5}$, and $I_{8}$ are traversed. Consider $\vec{x}_{4}$ as an example. From $I_{1}$, we obtain $x_{4}^{1}=6.8$, and the product $q^{1}\times x_{4}^{1}=17.0$ (this multiplication can be SIMD-accelerated) is temporarily stored in $T[0]$. Accumulating $T[0]$ into $A[4]$ gives $A[4]=0+17.0=17.0$. Similarly, from $I_{5}$ we compute $x_{4}^{5}\times q^{5}=14.0$, store it in $T[2]$, and add it to obtain $A[4]=31.0$. The computation for $I_{8}$ proceeds analogously. Finally, $A[4]$ becomes $36.1$, which equals $\delta(\vec{x}_{4},\vec{q})$. Although $\vec{x}_{4}$ has another non-zero entry $x_{4}^{2}$, it is not in $\Omega(\vec{x}_{4},\vec{q})$ and thus does not contribute to the inner product. The same accumulation process applies to $\vec{x}_{2}$, $\vec{x}_{3}$, $\vec{x}_{6}$, and $\vec{x}_{7}$. Eventually, $A[4]$ holds the largest value, so the nearest neighbor of $\vec{q}$ is $\vec{x}_{4}$.
Figure 4. An Example of Sindi and Query Process.
Using SIMD instructions, Sindi processes the $j$-th inverted list $I_{j}$ in batches, multiplying $q^{j}$ with each $x_{i}^{j}$ it contains and writing the results sequentially into $T^{j}$. This approach maximizes CPU utilization and reduces the complexity of the inner product computation from $\mathcal{O}(\lVert q\rVert+\lVert x\rVert)$ to $\mathcal{O}\!\left(\frac{\lVert q\rVert}{s}\right)$, where $s$ denotes the number of elements processed per SIMD operation. The complexity of Sindi’s distance computation is derived as follows:
@@
Let $\mathcal{D}$ be the dataset, $I$ its inverted index, and $I_{j}$ the posting list for dimension $j$. Let $\mathcal{I}^{j}=\{x_{i}^{j}\mid x_{i}^{j}\neq 0\}$ denote the set of non-zero entries in $I_{j}$.
Given a query vector $\vec{q}$, let $\mathcal{J}=\{\,j\mid q^{j}\neq 0\,\}$ be the set of query dimensions with non-zero entries. Let $\mathcal{X}=\{\,\vec{x}_{i}\mid\exists j\in\mathcal{J},\;x_{i}^{j}\neq 0\,\}$ denote the set of candidate vectors retrieved by $\vec{q}$.
Let $s$ be the number of elements that can be processed simultaneously using SIMD instructions. Then the amortized per-vector time complexity of computing the inner product between $\vec{q}$ and all $\vec{x}_{i}\in\mathcal{X}$ is $\Theta\!\left(\frac{\lVert q\rVert}{s}\right)$.
Proof. The total number of non-zero entries accessed in all posting lists for $\mathcal{J}$ is
$\sum_{j\in\mathcal{J}}\lVert\mathcal{I}^{j}\rVert.$
Since $s$ entries can be processed in parallel using SIMD, the total time is
$T_{\text{total}}=\sum_{j\in\mathcal{J}}\frac{\lVert\mathcal{I}^{j}\rVert}{s}.$
Amortizing over all $\lVert\mathcal{X}\rVert$ candidates gives
$T=\frac{T_{\text{total}}}{\lVert\mathcal{X}\rVert}=\frac{\sum_{j\in\mathcal{J}}\lVert\mathcal{I}^{j}\rVert}{s\cdot\lVert\mathcal{X}\rVert}.$
For any $\vec{x}_{i}\in\mathcal{X}$, we have $q\cap x_{i}\subseteq q$, implying that $\|\Omega(\vec{x}_{i},\vec{q})\|\leq\|q\|$. Therefore:
@@
Substituting into $T$ yields
$T\leq\frac{\sum_{\vec{x}_{i}\in\mathcal{X}}\lVert q\rVert}{s\cdot\lVert\mathcal{X}\rVert}=\frac{\lVert q\rVert}{s}.$
Hence, the amortized per-vector complexity is
$\Theta\!\left(\frac{\lVert q\rVert}{s}\right).$
@@
In summary, the Sindi index removes the overhead of redundant dimension matching and exploits SIMD parallelism to compute inner products in batches, thereby fully utilizing CPU computational resources. It also eliminates the memory access costs of traversing the original vectors, as the required values are retrieved directly from the posting lists and used for in-place product computation during retrieval.